OpenStudy (twizttiez):

Bob’s Boat Rentals charges a flat fee of $80 plus $5/h to rent a boat. Larry’s Boat Rentals charges $15/h with no flat fee. The system that models this situation is given, where c is the cost of renting a boat and h is the number of hours a boat is rented. The solution to the system is (8, 120). What is the correct interpretation of the solution of this system? c=80+5h c=15h

OpenStudy (twizttiez):

A. The cost to rent a boat is the same with both boat rental companies if you rent a boat for 8 h. Both companies will charge $120. B. The cost to rent a boat is the same with both boat rental companies if you rent a boat for 120 h. Both companies will charge $15. C. Bob’s Boat Rental will charge $120 for a rental of 8 hours. D. Bob’s Boat Rental will charge less money on the fifth day, charging $140.

OpenStudy (twizttiez):

@mathmate

OpenStudy (mathmate):

You have two equations representing the charges c related to h hours. Can you tell me what the solution means to you?

OpenStudy (twizttiez):

Solution = Answer

OpenStudy (twizttiez):

C (bob) = (flat fee) + (rate)*(hours) C (larry) = (rate)*(hours)

OpenStudy (mathmate):

Good, so far.

OpenStudy (mathmate):

 Here's a graphical representation of the problem that should be "flashing" in your mind.

OpenStudy (twizttiez):

Ok proceed

OpenStudy (mathmate):

Sorry, it's your turn to proceed, to tell me what is the meaning of the "solution" at (8,120). Does Larry charge more, or does Bob charge more for the 8 hour rental?

OpenStudy (twizttiez):

Larry charges more

OpenStudy (twizttiez):

c=80+h c=15h so 80 + h = 15h

OpenStudy (twizttiez):

subtract 1h from both sides then divide by 15 :)

OpenStudy (twizttiez):

then divide by 14

OpenStudy (twizttiez):

5.8

OpenStudy (twizttiez):

But after that what do i do

OpenStudy (mathmate):

Hmmm...??? C (bob) = (flat fee) + (rate)*(hours) = 80+5*8=? C (larry) = (rate)*(hours)=15*8 = ?

OpenStudy (twizttiez):

80+5*8=120 15*8 = 120

OpenStudy (twizttiez):

So it's A?

OpenStudy (mathmate):

Exactly!

OpenStudy (twizttiez):

Yay thanks so much have a great day!

OpenStudy (mathmate):

Note that C is also a true statement, but it is not the "appropriate" interpretation.

OpenStudy (twizttiez):

I was about to say C to lol

OpenStudy (mathmate):

You too! :)

OpenStudy (mathmate):

Exactly why I added the comment! :)